Grade 12Magical Mathematics[Interesting Approach]

the area oa a rectangle whose length is five more than twice its width is 75 square unit. the length is :- 5 unit OR 10 unit OR 15 unit OR 20 unit ?

askiit
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Let the width of the rectangle = x units
Length = (2 x + 5) units
According to the question,
Area = x(2x + 5)
=> 75 = 2x2+5x2x2+5x
=> 2x2+5x−75=02x2+5x−75=0
=> 2x2+15x−10x−75=02x2+15x−10x−75=0
=> x(2x+15)-5(2x+15)=0
=> (2x+15)(x-5)=0
=> x = 5 and −152−152
Width cannot be negative.
Width = 5 units
Length=2x+5 =2×5+52×5+5=15 units